OpenStudy (anonymous):

If g(x)=2x + 1, find g(a + h)- g(a).

OpenStudy (anonymous):

plz help

OpenStudy (jdoe0001):

@kingreggie bear in mind that \( g(x)=2x+1\\ \boldsymbol{g(\alpha+\theta)} \text{ it really means } \boldsymbol{x=\alpha+\theta}\\ \boldsymbol{g(\alpha)} \text{ it really means } \boldsymbol{x=\alpha} \)